OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (anonymous):

I got the second part, f(x) = ?(x-2)^2+6 I just need the first part.

OpenStudy (anonymous):

Answers: f(x) = −one eighth (x − 2)2 + 6 f(x) = one eighth (x − 2)2 + 6 f(x) = −one eighth (x + 2)2 + 8 f(x) = one eighth (x + 2)2 + 8

OpenStudy (anonymous):

never mind.

OpenStudy (anonymous):

The answer was A.